# Generating Physically Stable and Buildable LEGO® Designs from Text

### v1

@@ Abstract
We introduce LegoGPT, the first approach for generating physically stable LEGO brick models from text prompts. To achieve this, we construct a large-scale, physically stable dataset of LEGO designs, along with their associated captions, and train an autoregressive large language model to predict the next brick to add via next-token prediction. To improve the stability of the resulting designs, we employ an efficient validity check and physics-aware rollback during autoregressive inference, which prunes infeasible token predictions using physics laws and assembly constraints. Our experiments show that LegoGPT produces stable, diverse, and aesthetically pleasing LEGO designs that align closely with the input text prompts. We also develop a text-based LEGO texturing method to generate colored and textured designs. We show that our designs can be assembled manually by humans and automatically by robotic arms. We also release our new dataset, StableText2Lego, containing over 47,000 LEGO structures of over 28,000 unique 3D objects accompanied by detailed captions, along with our code and models at the project website: https://avalovelace1.github.io/LegoGPT/.
@@
Figure 1: Overview of LegoGPT. (a) Our method generates physically stable LEGO structures from text descriptions through an end-to-end approach, showing intermediate brick-by-brick steps. (b) The generated designs are buildable both by hand and by automated robotic assembly. (c) We show example results with corresponding text prompts. Besides basic LEGO designs (top), our method can generate colored LEGO models (bottom right) and textured models (bottom left) with appearance descriptions. We highly recommend the reader to check our website for step-by-step videos.
@@ 1 Introduction
3D generative models have made remarkable progress, driven by advances in generative modeling [15, 65] and neural rendering [26, 52]. These models have enabled various applications in virtual reality, gaming, entertainment, and scientific computing. For example, several works have explored synthesizing 3D objects from text [58], adding texture to meshes [61, 9], and manipulating the shape and appearance of existing 3D objects and scenes [40, 19].
@@
However, applying existing methods to create real-world objects remains challenging. Most approaches focus on generating diverse 3D objects with high-fidelity geometry and appearance [85, 21], but these digital designs often cannot be physically realized due to two key challenges [46]. First, the objects may be difficult to assemble or fabricate using standard components. Second, the resulting structure may be physically unstable even if assembly is possible. Without proper support, parts of the design can collapse, float, or remain disconnected.
@@
In this work, we address the challenge of generating physically realizable objects and study this problem in the context of LEGO® design. LEGO is widely used in entertainment, education, artistic creation, and manufacturing prototyping. Additionally, it can serve as a reproducible research benchmark, as all standard components are readily available. Due to the significant effort required for manual design, recent studies have developed automated algorithms to streamline the process and generate compelling results. However, existing approaches primarily create LEGO designs from a given 3D object [45] or focus on a single object category [13, 12].
@@
Our goal is to develop a method for generating LEGO designs directly from freeform text prompts while ensuring physical stability and buildability. Specifically, we aim to train a generative model that produces designs that are:
@@
  • Physically stable: Built on a LEGO baseplate with strong structural integrity, without floating or collapsing bricks.
@@
  • Buildable: Compatible with standard LEGO pieces and able to be assembled brick-by-brick by humans or robots.
@@
In this work, we introduce LegoGPT with the key insight of repurposing autoregressive large language models, originally trained for next-token prediction, for next-brick prediction. We formulate the problem of LEGO design as an autoregressive text generation task, where the next-brick dimension and placement are specified with a simple textual format. To ensure generated structures are both stable and buildable, we enforce physics-aware assembly constraints during both training and inference. During training, we construct a large-scale dataset of physically stable LEGO designs paired with captions. During autoregressive inference, we enforce feasibility with an efficient validity check and physics-aware rollback to ensure that the final tokens adhere to physics laws and assembly constraints.
@@
Our experiments show that the generated designs are stable, diverse, and visually appealing while adhering to input text prompts. Our method outperforms pre-trained LLMs with and without in-context learning, and previous approaches based on mesh-based 3D generation. Finally, we explore applications such as text-driven LEGO texturing, as well as manual assembly and automated robotic assembly of our designs. Our dataset, code, and models are available at the project website: https://avalovelace1.github.io/LegoGPT/.
@@ 2 Related Work
Text-to-3D generation has seen remarkable progress in recent years, driven by advances in neural rendering and generative models. Dreamfusion [58] and Score Jacobian Chaining [76] pioneer zero-shot text-to-3D generation by optimizing neural radiance fields [52] with pre-trained diffusion models [62]. Subsequent work has explored alternative 3D representations [35, 3, 48, 66, 33, 42, 51] and improved loss functions [77, 83, 47, 74, 25, 44]. Rather than relying on iterative optimization, a promising alternative direction trains generative models directly on 3D asset datasets, with various backbones including diffusion models [63, 89, 32, 54, 85, 34, 20, 86, 60], large reconstruction models [21, 31, 81, 73], U-Nets [68, 36], and autoregressive models [55, 4, 67, 64, 5, 6, 18, 79].
@@
However, these existing methods cannot be directly applied to LEGO generation because they do not account for the unique physical constraints and assembly requirements of real-world object designs [46]. Our work bridges this gap by introducing a framework for generating physically stable and buildable LEGO designs directly from text prompts.
Autoregressive 3D Modeling.
@@
Recent research has successfully used autoregressive models to generate 3D meshes [55, 64, 4, 67, 5, 6, 18, 79], often conditioned on input text or images. Most recently, LLaMA-Mesh [78] demonstrates that large language models (LLMs) can be fine-tuned to output 3D shapes in plain-text format, given a text prompt. However, most existing autoregressive methods focus on mesh generation. In contrast, we focus on generating LEGO designs from text prompts, leveraging LLMs’ reasoning capabilities.
@@
Figure 2: StableText2Lego Dataset. (a) From a ShapeNetCore [2] mesh, we generate a LEGO design by voxelizing it onto a $20\times 20\times 20$ grid and applying legolization to determine the brick layout. (b) We augment each shape with multiple structural variations by randomizing the brick layout while preserving the overall shape. (c) Stability analysis [37] is performed on each variation to filter out physically unstable designs. (d) To obtain the corresponding captions for each shape, we render the LEGO design from 24 different viewpoints and use GPT-4o [1] to generate detailed geometric descriptions. (e) Data samples from 5 categories in our StableText2Lego dataset.
LEGO Assembly and Generation.
@@
Creating LEGO® designs given a reference 3D shape has been widely studied [28]. Existing works [57, 69, 88] formulate the generation as an optimization problem guided by hand-crafted heuristic rules. Such heuristics can include ensuring that all bricks are interconnected, minimizing the number of bricks, and maximizing the number of brick orientation alternations. Luo et al. [45] leverages structural stability estimation to find weak structural parts and rearrange the local brick layout to generate physically stable designs. Kim et al. [27], Liu et al. [39] formulate a planning problem to fill the target 3D model sequentially. However, these methods only generate designs given an input 3D shape, assuming a valid LEGO design exists, which is difficult to verify in practice.
@@
Few works have explored learning-based techniques to generate LEGO designs. Thompson et al. [70] use a deep graph generative model in which the graph encodes brick connectivity. However, this method is limited to generating simple classes, such as walls and cuboids, using a single brick type. More recently, Ge et al. [13] use a diffusion model to predict a semantic volume, which is then translated into a high-quality micro building. Their method produces impressive results for a single category. Zhou et al. [87] and Ge et al. [12] generate compelling figurine designs given an input portrait. They use machine learning to select from a pre-made set of components that best match an input photo. Although effective for faces, extending this selection-based approach to arbitrary objects is challenging. Zhou et al. [90] formulate an optimization problem to create a LEGO model from an input image. While their output is a 2D LEGO mosaic, we focus on 3D structures in this work. Goldberg et al. [14] queries a vision-language model (VLM) to generate diverse 3D assembly structures. However, they use regular building blocks instead of LEGO bricks, which do not have interlocking connections and thus have limited expressiveness.
@@
Our goal is closest to that of [30]. This work has three steps: (1) generating an image using a text-to-image model, (2) converting the image into voxels, and (3) using heuristics to create a physical LEGO build without considering physical constraints. In contrast, our method performs the text-to-LEGO task without requiring intermediate image or voxel representations.
Physics-Aware Generation.
@@
Despite significant advances in 3D generation, incorporating physics constraints remains largely underexplored; most approaches focus primarily on geometry or appearance modeling. Physics-aware generation can be broadly categorized into two approaches: direct constraint enforcement and learned validation. Simple physical constraints, such as collision avoidance and contact requirements, can be incorporated directly through explicit penalty terms during optimization [53, 82, 84, 75, 41, 24, 16]. More complex physical properties, such as structural stability and dynamic behavior, typically require differentiable physics simulators [7, 56, 80, 50] or data-driven physics-aware assessment models [10, 49]. To our knowledge, our paper is the first attempt to incorporate physics-aware constraints into text-based LEGO generation.
@@ 3 Dataset
Training a modern autoregressive model requires a large-scale dataset. Therefore, we introduce StableText2Lego, a new large-scale LEGO® dataset that contains $47{,}000+$ different LEGO structures, covering $28{,}000+$ unique 3D objects from $21$ common object categories of the ShapeNetCore dataset [2]. We select categories featuring diverse and distinctive 3D objects while excluding those resembling cuboids. Each structure is paired with a group of text descriptions and a stability score, which indicates its physical stability and buildability. Below, we describe the dataset construction, an overview of which is given in Figure 2.
@@
Figure 3: Method. (a) Our system tokenizes a LEGO design into a sequence of text tokens, ordered in a raster-scan manner from bottom to top. (b) We create an instruction dataset pairing brick sequences with descriptions to fine-tune LLaMA-3.2-Instruct-1B. (c) At inference time, LegoGPT generates LEGO designs incrementally by predicting one brick at a time given a text prompt. For each generated brick, we perform validity checks to ensure it is well-formatted, exists in our brick library, and does not collide with existing bricks. After completing the design, we verify its physical stability. If the structure is unstable, we roll back to the stable state by removing all the unstable bricks and their subsequent ones and resume generation from that point.
LEGO Representation.
@@
We consider LEGO structures built on a LEGO baseplate. Each LEGO structure in StableText2Lego is represented as $B=[b_{1},b_{2},\dots,b_{N}]$ with $N$ bricks, and each element denotes a brick’s state as $b_{i}=[h_{i},w_{i},x_{i},y_{i},z_{i}]$. Here, $h_{i}$ and $w_{i}$ indicate the brick length in the $X$ and $Y$ directions, respectively, and $x_{i}$, $y_{i}$, and $z_{i}$ denote the position of the stud closest to the origin. The position has $x_{i}\in[0,1,\dots,H-1]$, $y_{i}\in[0,1,\dots,W-1]$, $z_{i}\in[0,1,\dots,D-1]$, and $H$, $W$, and $D$ represent the dimensions of the discretized grid world.
Mesh-to-LEGO.
@@
We construct the dataset by converting 3D shapes from ShapeNetCore [2] into LEGO structures as shown in Figure 2(a). Given a 3D mesh, we voxelize and downsample it to a $20\times 20\times 20$ grid world to ensure a consistent scale, i.e., $H=W=D=20$. The LEGO brick layout is generated by a split-and-remerge legolization algorithm. To improve data quality and diversity, we introduce randomness in legolization and generate multiple different LEGO structures for the same 3D object, as illustrated in Figure 2(b). This increases the chance of obtaining a stable LEGO structure and more diverse LEGO brick layouts. More details can be found in Appendix A. We use eight commonly available standard bricks: $1\times 1$, $1\times 2$, $1\times 4$, $1\times 6$, $1\times 8$, $2\times 2$, $2\times 4$, and $2\times 6$.
Stability Score.
@@
We assess the physical stability of each structure, as illustrated in Figure 2(c), using the analysis method [37]. For a structure $B=[b_{1},b_{2},\dots,b_{N}]$, the stability score $S\in\mathbb{R}^{N}$ assigns each brick $b_{i}$ a value $s_{i}\in[0,1]$ that quantifies the internal stress at its connections. Higher scores $s_{i}$ indicate greater stability, while $s_{i}=0$ indicates an unstable brick that will cause structural failure. Calculating the stability score requires solving a nonlinear program to determine the forces acting on each brick to achieve a static equilibrium that prevents structural collapse, as detailed in Section 4.2. For typically-sized (i.e., $<200$ bricks) structures in Figure 2, it takes $\sim$0.35 seconds on average. We only include stable structures where all bricks have stability scores greater than 0.
Caption Generation.
@@
To obtain captions for each shape, we render the LEGO from 24 different viewpoints and combine them into a single multi-view image. We then prompt GPT-4o [1] to produce five descriptions for these renderings with various levels of detail. Importantly, we ask GPT-4o to omit color information and focus only on geometry. The complete GPT-4o prompt is provided in Appendix A.
@@
Figure 2(e) shows several data samples in StableText2Lego. The rich variations within each category and the comprehensive text-LEGO pairs make it possible to train large-scale generative models. More insights on our StableText2Lego are discussed in Appendix A.

### v2

# Generating Physically Stable and Buildable Brick Structures from Text

@@ Abstract
We introduce BrickGPT, the first approach for generating physically stable interconnecting brick assembly models from text prompts. To achieve this, we construct a large-scale, physically stable dataset of brick structures, along with their associated captions, and train an autoregressive large language model to predict the next brick to add via next-token prediction. To improve the stability of the resulting designs, we employ an efficient validity check and physics-aware rollback during autoregressive inference, which prunes infeasible token predictions using physics laws and assembly constraints. Our experiments show that BrickGPT produces stable, diverse, and aesthetically pleasing brick structures that align closely with the input text prompts. We also develop a text-based brick texturing method to generate colored and textured designs. We show that our designs can be assembled manually by humans and automatically by robotic arms. We release our new dataset, StableText2Brick, containing over 47,000 brick structures of over 28,000 unique 3D objects accompanied by detailed captions, along with our code and models at the project website: https://avalovelace1.github.io/BrickGPT/.
@@
Figure 1: Overview of BrickGPT. (a) Our method generates physically stable interconnecting brick assembly structures from text descriptions through an end-to-end approach, showing intermediate brick-by-brick steps. (b) The generated designs are buildable both by hand and by automated robotic assembly. (c) We show example results with corresponding text prompts. Besides basic brick designs (top), our method can generate colored brick models (bottom right) and textured models (bottom left) with appearance descriptions. We highly recommend the reader to check our website for step-by-step videos.
@@ 1 Introduction
3D generative models have made remarkable progress, driven by advances in generative modeling [16, 67] and neural rendering [27, 53]. These models have enabled various applications in virtual reality, gaming, entertainment, and scientific computing. Several works have explored synthesizing 3D objects from text [60], adding texture to meshes [63, 8], and manipulating the shape and appearance of existing 3D objects and scenes [41, 20].
@@
However, creating real-world objects with existing methods remains challenging. Most approaches focus on generating diverse 3D objects with high-fidelity geometry and appearance [88, 22], but these digital designs often cannot be physically realized due to two key challenges [47]. First, the objects may be difficult to assemble or fabricate using standard components. Second, the resulting structure may be physically unstable even if assembly is possible. Without proper support, parts of the design could collapse, float, or remain disconnected.
@@
In this work, we address the challenge of generating physically realizable objects. We study this problem in the context of designing structures made of interlocking toy bricks, such as LEGO® blocks. These are widely used in entertainment, education, artistic creation, and manufacturing prototyping. Additionally, they can serve as a reproducible research benchmark, as all standard components are readily available. Due to the significant effort required to design brick structures manually, recent studies have developed automated algorithms to streamline the process and generate compelling results. However, existing approaches primarily create structures from a given 3D object [46] or focus on a single object category [14, 13].
@@
Our goal is to develop a method for generating brick assembly structures directly from freeform text prompts while ensuring physical stability and buildability. Specifically, we aim to train a generative model that produces designs that are:
@@
  • Physically stable: Built on a baseplate with strong structural integrity, without floating or collapsing bricks.
@@
  • Buildable: Compatible with standard interconnecting toy brick pieces and able to be assembled brick-by-brick by humans or robots.
@@
In this work, we introduce BrickGPT with the key insight of repurposing autoregressive large language models, originally trained for next-token prediction, for next-brick prediction. We formulate the problem of brick structure design as an autoregressive text generation task, where the next-brick dimension and placement are specified with a simple textual format.
To ensure generated structures are both stable and buildable, we enforce physics-aware assembly constraints during both training and inference. During training, we construct a large-scale dataset of physically stable brick structures paired with captions. During autoregressive inference, we enforce feasibility with an efficient validity check and physics-aware rollback to ensure that the final tokens adhere to physics laws and assembly constraints.
@@
Our experiments show that the generated designs are stable, diverse, and visually appealing while adhering to input text prompts. Our method outperforms pre-trained LLMs with and without in-context learning, and previous approaches based on 3D mesh generation. Finally, we explore applications such as text-driven brick texturing, as well as manual assembly and automated robotic assembly of our designs. Our dataset, code, and models are available at the project website: https://avalovelace1.github.io/BrickGPT/.
@@ 2 Related Work
Text-to-3D Generation.   Text-to-3D generation has seen remarkable progress in recent years, driven by advances in neural rendering and generative models. Dreamfusion [60] and Score Jacobian Chaining [78] pioneer zero-shot text-to-3D generation by optimizing neural radiance fields [53] with pre-trained diffusion models [64]. Subsequent work has explored alternative 3D representations [36, 3, 49, 68, 34, 43, 52] and improved loss functions [80, 86, 48, 76, 26, 45]. Rather than relying on iterative optimization, a promising alternative direction trains generative models directly on 3D asset datasets, with various backbones including diffusion models [65, 92, 33, 55, 88, 35, 21, 89, 62], large reconstruction models [22, 32, 84, 75], U-Nets [70, 37], and autoregressive models [56, 4, 69, 66, 5, 6, 19, 82].
@@
However, these existing methods cannot be directly applied to generating brick structures because they do not account for the unique physical constraints and assembly requirements of real-world designs [47]. Our work bridges this gap by introducing a method for generating physically stable and buildable brick structures directly from text prompts.
@@
Autoregressive 3D Modeling.   Recent research has successfully used autoregressive models to generate 3D meshes [56, 66, 4, 69, 5, 6, 19, 82, 9], often conditioned on input text or images. Most recently, LLaMA-Mesh [81] demonstrates that large language models (LLMs) can be fine-tuned to output 3D shapes in plain-text format, given a text prompt. However, most existing autoregressive methods focus on mesh generation. In contrast, we focus on generating brick structures from text prompts, leveraging LLMs’ reasoning capabilities.
@@
Figure 2: StableText2Brick Dataset. (a) From a ShapeNetCore [2] mesh, we generate a brick structure by voxelizing it onto a $20\times 20\times 20$ grid, then constructing its brick layout with a delete-and-rebuild algorithm. (b) We augment each shape with multiple structural variations by randomizing the brick layout while preserving the overall shape. (c) Stability analysis [38] is performed on each variation to filter out physically unstable designs. (d) To obtain captions for each shape, we render the brick structure from 24 different viewpoints and use GPT-4o [1] to generate detailed geometric descriptions. (e) Data samples from 5 categories in our StableText2Brick dataset.
@@
Brick Assembly and Design Generation.   Creating brick structures given a reference 3D shape has been widely studied [29]. Existing works [58, 71, 91] formulate the generation as an optimization problem guided by hand-crafted heuristic rules. Such heuristics can include ensuring that all bricks are interconnected, minimizing the number of bricks, and maximizing the number of brick orientation alternations. Wang et al. [79] translate a visual manual into step-by-step brick assembly instructions. Luo et al. [46] leverage structural stability estimation to find weak structural parts and rearrange the local brick layout to generate physically stable designs. Kim et al. [28], Liu et al. [40] formulate a planning problem to fill the target 3D model sequentially. However, these methods only generate designs given an input 3D shape, assuming a valid brick structure exists, which is difficult to verify in practice.
@@
Few works have explored learning-based techniques to generate toy brick designs. Thompson et al. [72] use a deep graph generative model in which the graph encodes brick connectivity. However, this method is limited to generating simple classes, such as walls and cuboids, using a single brick type. More recently, Ge et al. [14] use a diffusion model to predict a semantic volume, which is then translated into a high-quality micro building. Their method produces impressive results for a single category. Zhou et al. [90] and Ge et al. [13] generate compelling figurine designs given an input portrait. They use machine learning to select from a pre-made set of components that best match an input photo. Although effective for faces, extending this selection-based approach to arbitrary objects is challenging. Zhou et al. [93] formulate an optimization problem to create a brick model from an input image. While their output is a 2D brick mosaic, we focus on 3D structures in this work. Goldberg et al. [15] query a vision-language model to generate diverse 3D assembly structures. However, they use regular building blocks instead of bricks with interlocking connections, and thus the structures have limited expressiveness.
@@
Our goal is closest to that of [31]. This work has three steps: (1) generating an image using a text-to-image model, (2) converting the image into voxels, and (3) using heuristics to create a physical brick model without considering physical constraints. In contrast, our method performs the text-to-brick task without requiring intermediate image or voxel representations.
@@
Physics-Aware Generation.   Physics-aware 3D generation can be broadly categorized into two approaches: direct constraint enforcement and learned validation. Simple physical constraints, such as collision avoidance and contact requirements, can be incorporated directly through explicit penalty terms during optimization [54, 85, 87, 77, 42, 25, 17]. More complex physical properties, such as structural stability and dynamic behavior, typically require physics simulators [12, 57, 83, 51] or data-driven physics-aware assessment models [10, 50]. To our knowledge, our paper is the first attempt to incorporate physics-aware constraints into text-based brick assembly structure generation.
@@ 3 Dataset
Training a modern autoregressive model requires a large-scale dataset. Therefore, we introduce StableText2Brick, a new large-scale dataset that contains $47{,}000+$ different toy brick assembly structures, covering $28{,}000+$ unique 3D objects from $21$ common object categories of the ShapeNetCore dataset [2]. We select categories featuring diverse and distinctive 3D objects while excluding those resembling cuboids. Each structure is paired with a group of text descriptions and a stability score, which indicates its physical stability and buildability. Below, we describe the dataset construction, an overview of which is given in Figure 2.
@@
Figure 3: Method. (a) Our system tokenizes a brick structure into a sequence of text tokens, ordered in a raster-scan manner from bottom to top. (b) We create an instruction dataset pairing brick sequences with descriptions to fine-tune LLaMA-3.2-Instruct-1B. (c) At inference time, BrickGPT generates brick structures incrementally by predicting one brick at a time given a text prompt. For each generated brick, we perform validity checks to ensure it is well-formatted, exists in our brick library, and does not collide with existing bricks. After completing the design, we verify its physical stability. If the structure is unstable, we roll back to a stable state by removing all unstable bricks and their subsequent bricks, and resume generation from that point.
@@
Brick Representation.   We consider brick structures built on a fixed baseplate. Each structure in StableText2Brick is represented as $B=[b_{1},b_{2},\dots,b_{N}]$ with $N$ bricks, and each element denotes a brick’s state as $b_{i}=[h_{i},w_{i},x_{i},y_{i},z_{i}]$. Here, $h_{i}$ and $w_{i}$ indicate the brick length in the $X$ and $Y$ directions, respectively, and $x_{i}$, $y_{i}$, and $z_{i}$ denote the position of the stud closest to the origin. The position has $x_{i}\in[0,1,\dots,H-1]$, $y_{i}\in[0,1,\dots,W-1]$, $z_{i}\in[0,1,\dots,D-1]$, where $H$, $W$, and $D$ represent the dimensions of the discretized grid world.
@@
Mesh-to-Brick.   We construct the dataset by converting 3D shapes from ShapeNetCore [2] into brick structures as shown in Figure 2(a). Given a 3D mesh, we voxelize and downsample it into a $20\times 20\times 20$ grid world to ensure a consistent scale, i.e., $H=W=D=20$. The brick layout is generated by a delete-and-rebuild algorithm similar to [46]. To improve data quality and diversity, we introduce randomness and generate multiple different structures for the same 3D object, as illustrated in Figure 2(b). This increases the chance of obtaining a stable structure and more diverse layouts. More details can be found in Appendix A. We use eight commonly available standard bricks: $1\times 1$, $1\times 2$, $1\times 4$, $1\times 6$, $1\times 8$, $2\times 2$, $2\times 4$, and $2\times 6$.
@@
Stability Score.   We assess the physical stability of each structure, as illustrated in Figure 2(c), using the analysis method [38]. For a structure $B=[b_{1},b_{2},\dots,b_{N}]$, the stability score $S\in\mathbb{R}^{N}$ assigns each brick $b_{i}$ a value $s_{i}\in[0,1]$ that quantifies the internal stress at its connections. Higher scores $s_{i}$ indicate greater stability, while $s_{i}=0$ indicates an unstable brick that will cause structural failure. Calculating the stability score requires solving a nonlinear program to determine the forces acting on each brick to achieve a static equilibrium that prevents structural collapse, as detailed in Section 4.2. For typically-sized (i.e., $<200$ bricks) structures in Figure 2, stability analysis takes $\sim$0.35 seconds on average. A structure is stable if all bricks have stability scores greater than 0; we only include stable structures in the StableText2Brick dataset.
@@
Caption Generation.   To obtain captions for each shape, we render the brick structure from 24 different viewpoints and combine them into a single multi-view image. We then prompt GPT-4o [1] to produce five descriptions for these renderings with various levels of detail. Importantly, we ask GPT-4o to omit color information and focus only on geometry. The complete prompt is provided in Appendix A.
@@
Figure 2(e) shows several data samples in StableText2Brick. The rich variations within each category and the comprehensive text-brick pairs make it possible to train large-scale generative models. More insights on StableText2Brick are discussed in Appendix A.